OpenStudy (anonymous):

Find the general solution of y'=x*sqrt(x+1)

OpenStudy (kc_kennylau):

I think it's integral-by-part

OpenStudy (anonymous):

are you trying to solve for x? or looking for the integral?

OpenStudy (anonymous):

u=x+1 du=dx

OpenStudy (kc_kennylau):

He's solving for y

OpenStudy (anonymous):

if he is solving for y the answer stays the same. (I think) unless he wants the inverse of y' which will be \[y^2=x^3+x^2\]

OpenStudy (anonymous):

in that case y=\[y=-\frac{ 10(x+1)^3/2}{ 15 }+\frac{ 6(x+1)^5/2 }{ 15 }\]